Find the integral of (y2 -1) dx-2dy = 0


1
Expert's answer
2022-04-11T12:09:40-0400

"(y^2-1)dx=2dy,"

"\frac 12 x=\frac 12\ln\frac{1-y}{1+y}+\frac 12 \ln c,"

"y=\frac{1-e^{2c+x}}{1+e^{2c+x}}."